# Prompt-driven Transferable Adversarial Attack on Person Re-Identification with Attribute-aware Textual Inversion

### v2

@@ Abstract
Person re-identification (re-id) models are vital in security surveillance systems, requiring transferable adversarial attacks to explore the vulnerabilities of them. Recently, vision-language models (VLM) based attacks have shown superior transferability by attacking generalized image and textual features of VLM, but they lack comprehensive feature disruption due to the overemphasis on discriminative semantics in integral representation. In this paper, we introduce the Attribute-aware Prompt Attack (AP-Attack), a novel method that leverages VLM’s image-text alignment capability to explicitly disrupt fine-grained semantic features of pedestrian images by destroying attribute-specific textual embeddings. To obtain personalized textual descriptions for individual attributes, textual inversion networks are designed to map pedestrian images to pseudo tokens that represent semantic embeddings, trained in the contrastive learning manner with images and a predefined prompt template that explicitly describes the pedestrian attributes. Inverted benign and adversarial fine-grained textual semantics facilitate attacker in effectively conducting thorough disruptions, enhancing the transferability of adversarial examples. Extensive experiments show that AP-Attack achieves state-of-the-art transferability, significantly outperforming previous methods by 22.9% on mean Drop Rate in cross-model&dataset attack scenarios.
@@ 3.1 Preliminaries
0:  Batch images $x$, visual encoder $\mathcal{T}$ and textual encoder $\mathcal{V}$, prompt template $p$, surrogate model $\mathcal{M}$.
@@ 4.3 Ablation Studies
The effectiveness of textual inversion. In order to verify the effectiveness of our textual inversion networks, we attempt to interpret the learned pseudo-tokens to meaningful word. We first established an attribute-specific semantic vocabulary  [^1] The full details of produced semantic vocabulary shows in supplementary files. by chatGPT, with each attribute corresponding to a distinct, meaningful set of semantic words. These attribute vocabularies capture both color and descriptive features. Then, we calculate the similarity between each word in the vocabulary and the pseudo-tokens, selecting the two words with the highest similarity scores for display, as shown in Fig. 3. From the figure, it can be seen that the semantic words similar to the learned pseudo-tokens can correctly describe the pedestrian image information, indicating the effectiveness of our inversion networks.
@@ 4.4 Attack Effectiveness against Defense Method
+global visual feature 27.9 63.6
+integral text embedding 25.9 66.2
+fine-grained semantic embedding 21.2 72.4
@@ 4.5 Transferability to Diverse Types Models
To further evaluate our AP-Attack’s generalizability across diverse types of re-id models, we test it against three distinct re-id models: the self-supervised PASS (Market) [70], auxiliary-feature-enhanced PGFA (Occuluded-Duke) [31], and CLIP-ReID (Market) [22] based on CLIP. Tab. 6 reveals substantial performance degradation across all tested architectures, confirming the efficacy of our method across diverse model paradigms.
@@ 5 Conclusion
In this paper, we propose a novel Attribute-aware Prompt Attack methods to enhance the transferability of adversarial attacks on person re-id task. Our AP-Attack method leverages the image-text alignment capability of VLM and introduces the attribute-specific inversion networks to map the image feature to attribute semantic textual embeddings. And it attempts to thoroughly destroy the pedestrian features by perturbing fine-grained attribute semantics across all attribute feature spaces. Extensive experimental results validate the superiority of our approach in cross-dataset and cross-model black-box attack scenarios, achieving substantial performance gains over the latest SOTA methods. We believe that our work offers a meaningful contribution to adversarial attack research and holds promise for strengthening the security of machine learning systems in real-world applications.

### v3

@@ Abstract
Person re-identification (re-id) models are vital in security surveillance systems, requiring transferable adversarial attacks to explore the vulnerabilities of them. Recently, vision-language models (VLM) based attacks have shown superior transferability by attacking generalized image and textual features of VLM, but they lack comprehensive feature disruption due to the overemphasis on discriminative semantics in integral representation. In this paper, we introduce the Attribute-aware Prompt Attack (AP-Attack), a novel method that leverages VLM’s image-text alignment capability to explicitly disrupt fine-grained semantic features of pedestrian images by destroying attribute-specific textual embeddings. To obtain personalized textual descriptions for individual attributes, textual inversion networks are designed to map pedestrian images to pseudo tokens that represent semantic embeddings, trained in the contrastive learning manner with images and a predefined prompt template that explicitly describes the pedestrian attributes. Inverted benign and adversarial fine-grained textual semantics facilitate attacker in effectively conducting thorough disruptions, enhancing the transferability of adversarial examples. Extensive experiments show that AP-Attack achieves state-of-the-art transferability, significantly outperforming previous methods by 22.9% on mean Drop Rate in cross-model&dataset attack scenarios. The codes are at https://github.com/yuanbianGit/AP-Attack.
@@ 3.1 Preliminaries
0:  Batch images $x$, visual encoder $\mathcal{V}$ and textual encoder $\mathcal{T}$, prompt template $p$, surrogate model $\mathcal{M}$.
@@ 4.3 Ablation Studies
The effectiveness of textual inversion. In order to verify the effectiveness of our textual inversion networks, we attempt to interpret the learned pseudo-tokens to meaningful word. We first established an attribute-specific semantic vocabulary Details of produced semantic vocabulary shows in supplemental files. by chatGPT, with each attribute corresponding to a distinct, meaningful set of semantic words. These attribute vocabularies capture both color and descriptive features. Then, we calculate the similarity between each word in the vocabulary and the pseudo-tokens, selecting the two words with the highest similarity scores for display, as shown in Fig. 3. From the figure, it can be seen that the semantic words similar to the learned pseudo-tokens can correctly describe the pedestrian image information, indicating the effectiveness of our inversion networks.
@@ 4.4 Attack Effectiveness against Defense Method
baseline+global visual feature 27.9 63.6
baseline+integral text embedding 25.9 66.2
baseline+fine-grained semantic embedding 21.2 72.4
@@ 4.5 Transferability to Diverse Types Models
To further evaluate our AP-Attack’s generalizability across diverse types of re-id models, we test it against three distinct re-id models: the self-supervised PASS (Market) [70], auxiliary-feature-enhanced PGFA (Occluded-Duke) [31], and CLIP-ReID (Market) [22] based on CLIP. Tab. 6 reveals substantial performance degradation across all tested architectures, confirming the efficacy of our method across diverse model paradigms.
@@ 5 Conclusion
In this paper, we propose a novel Attribute-aware Prompt Attack methods to enhance the transferability of adversarial attacks on person re-id task. Our AP-Attack method leverages the image-text alignment capability of VLM and introduces the attribute-specific inversion networks to map the image feature to attribute semantic textual embeddings. And it attempts to thoroughly destroy the pedestrian features by perturbing fine-grained attribute semantics across all attribute feature spaces. Extensive experimental results validate the superiority of our approach in cross-dataset and cross-model black-box attack scenarios, achieving substantial performance gains over the latest SOTA methods. We believe that our work offers a meaningful contribution to adversarial attack research and holds promise for strengthening the security of machine learning systems in real-world.
Acknowledgments. This work was supported in part by the National Natural Science Foundation of China under Grant 62425305, U22B2050 and 62221002, in part by the Science and Technology Innovation Program of Hunan Province under Grant 2023RC1048, in part by the Hunan Provincial Natural Science Foundation of China under Grant 2024JJ3013.